# Valuation rings as limits of complete intersection rings

### v11

@@ 1. Filtered colimits of smooth algebras
The following lemma is useful in the next section to reduce the study of filtered colimit of complete intersection algebra to the case of complete valuation rings.
Lemma 20.
Let $V\subset V^{\prime}$ be an extension of valuation rings and $^{\hat{V}^{\prime}}$ the completion of $V^{\prime}$. Assume that the field extension of $V\subset V^{\prime}$ is separable. Then $V^{\prime}$ is a filtered colimit of complete intersection $V$-algebras if $^{\hat{V}}^{\prime}$ is a filtered colimit of complete intersection $V$-algebras. Moreover, if the above fraction field extension is not separable, $E$ is a finitely presented $V$-algebra and $w:E\rightarrow V^{\prime}$ is a $V$-morphism such that $w(H_{E/V})V^{\prime}\not=0$ then $w$ factors through a complete intersection $V$-algebra if $^{\hat{V}}^{\prime}$ is a filtered colimit of complete intersection $V$-algebras.
Proof.
We follow the proofs of [21, Lemma 7, Proposition 9]. Assume that $^{\hat{V}}^{\prime}$ is a filtered colimit of complete intersection $V$-algebras. Let $E$ be a finitely generated $V$-subalgebra of $V^{\prime}$ and $w$ the inclusion $\subset V^{\prime}. EE$ is finitely presented by [15, Theorem 4]. By separability of the fraction field extension of $V\subset E$, the singular locus $w(H_{E/V})V^{\prime}$ has the form $zV^{\prime}$ for some nonzero $z\in V^{\prime}$. The same thing follows when the separability is missing but $w(H_{E/V})V^{\prime}\not=0$ . Actually, changing if necessary $E$ by $E[Z]$ and $w$ by the possible non injective map $E[Z]\rightarrow V^{\prime}$ given by $Z\rightarrow z$ we may assume that $w(H_{E/V})V^{\prime}$ has the form $dV^{\prime}$ for some $d\in E$.
Suppose that $E\cong V[Y]/I$, $Y=(Y_{1}\ldots,Y_{m})$, $I$ being a finitely generated ideal. As in the proof of [21, Proposition 20] we may assume that for some polynomials $f=(f_{1},\ldots,f_{r})$ from $I$, we have $d\in NME$ for some $N\in((f):I)$ and a $r\times r$-minor $M$ of the Jacobian matrix $(\partial f_{i}/\partial Y_{j})$.
Note that $^{\hat{w}}$ the composite map $E\xrightarrow{w}V^{\prime}\subset{\hat{V}}^{\prime}$ factors through a complete intersection $V$-algebra $B$, let us say $\hat{w}$ is the composite map $E\rightarrow B\xrightarrow{\rho^{\prime}}{\hat{V}}^{\prime}$. We may assume that $B$ has the form $T_{h}/(P)$ for a polynomial $V$-algebra $T=V[U]$, $U=(U_{1},\ldots,U_{l})$, $h\in T$ and a regular system of elements $P=(P_{1},\ldots,P_{t})$ from $T$. Suppose the map $\rho^{\prime}$ is given by $U\rightarrow{\hat{u}}$ for some ${\hat{u}}\in({\hat{V}}^{\prime})^{l}$. Thus h(u^)h($\hat${u}}) is invertible. Clearly, the map $\bar{w}$ induced by $w$ modulo $d^{3}$ factors through ${\bar{B}}=B/(d^{3})$, that is $\bar{w}$ is the composite map ${\bar{E}}=E/(d^{3})\rightarrow{\bar{B}}\rightarrow{\bar{V}}^{\prime}=V^{\prime}/(d^{3})\cong{\hat{V}}^{\prime}/(d^{3})$, the last map being given by $U\rightarrow u+d^{3}({\hat{V}}^{\prime})^{l}$, for some $u\in(V^{\prime})^{l}$ which is $\hat{u}$ modulo $d^{3}$. We have $P(u)=d^{3}z$ for some $z\in V^{\prime t}$. Set $D=(V[U,Z]/(P-d^{3}Z))_{h}$, $Z=(Z_{1},\ldots,Z_{t})$ and let $\rho:D\rightarrow V^{\prime}$ be the map $(U,Z)\rightarrow(u,z)$. Note that $D$ is a complete intersection $V$-algebra. As $w$ factors through $\rho$ modulo $d^{3}$ there exists a smooth $D$-algebra $D^{\prime}$ such that $w$ factors through $D^{\prime}$ using Lemma 11 applied for $A=V$ and $A^{\prime}=D$. It follows that $D^{\prime}$ is a complete intersection $V$-algebra as well and it is enough to apply [28, Lemma 1.5].
@@ 2. Model Theory of valued fields
The conclusion goes from Corollary 7 if we show that $k\subset V$. A lifting of $k$ to $V$ could be done when $V$ is Henselian and char $k=0$ (see [31, Theorem 2.9] but the proof goes in the same way in positive characteristic and $k$ is separably generated over $F$. In particular the lifting could be done when $k/F$ is of finite type.
@@
where $\widetilde{K}$ and $\widetilde{\Gamma}$ are the fraction field and the value group of $\widetilde{V}$.
@@
Let $V$ be a valuation ring containing its residue field $k$ of positive characteristic and $\Gamma$ its value group. Then $V$ is a filtered colimit of complete intersection $k$-algebras.
@@
As $K_{E}$ is a finite type field extension of $F$ we may assume $\Gamma$ is countable generated and $k$ is countable generated over $F$ by [3, Corollary 5.2]. Consider $U_{i}$, $\mathcal{U}_{i}$, $\widetilde{V}$, $\widetilde{k}$, $\widetilde{\Gamma}$, $\widetilde{s}$, $\widetilde{t}$ given by Theorem 23 for $V$. Note that $\widetilde{k}\subset\widetilde{V}$ because $k\subset V$. As usual $\widetilde{V}$ is an immediate extension of a valuation ring $)\widetilde{V}_{0}=\widetilde{V}\cap\widetilde{k}(\widetilde{s}(\widetilde{\Gamma}$)), which is a filtered union of localizations of polynomial algebras over $\widetilde{k}$. Using Theorem 6 we see that $\widetilde{V}$ is a filtered union of its complete intersection $\widetilde{V}_{0}$-subalgebras.
Then the map $w^{\prime}:\widetilde{k}\otimes_{F}E\rightarrow\widetilde{V}$ factors through a complete intersection $\widetilde{k}-$algebra and so through a complete intersection $k$-algebra $D$ since $\widetilde{k}/k$ is separable. Thus $w$ factors through $D$ too because all finite systems of polynomial equations which have a solution in $\tilde{V}$ must have one in $V$. The proof ends applying [28, Lemma 1.5].
@@
Let $V$ be a valuation ring with value group $\Gamma$ and $\tau=card\ \Gamma$. Then there exists an ultrafilter $\mathcal{U}$ on a set $U$ with $card\ U=\tau$ such that any system of polynomial equations $(g_{i}((X_{j})_{j\in J})_{i\in I}$ with $card\ I\leq\tau$ in variables $(X_{j})_{j\in J}$ with coefficients in the ultrapower ${\tilde{V}}=\Pi_{\mathcal{U}}V$ has a solution in $\tilde{V}$} if and only if all its finite subsystems have.
@@
Let $V$ be a valuation ring containing a field $F$ of positive characteristic. Assume that the residue field $k$ of $V$ is a separable extension of $F$ and one of the following conditions holds
  1. (1)
    $V$ is Henselian,
  2. (2)
    $\dim V=1$ and $V$ is a separable extension of $F$.
Then $V$ is a filtered colimit of complete intersection $F$-algebras.
@@
Let $\Gamma$ be the value group of $V$. Let $E$ be a finitely generated $F$-subalgebra (in fact finitely presented by [15, Theorem 4]) of $V$ and $w:E\rightarrow V$ the inclusion. We will show that $w$ factors through a complete intersection $V$-algebra. It is enough to replace $V$ by the intersection of $V$ with the fraction field $K_{E}$ of $E$. Thus we may assume that $\Gamma$ is countable generated and $k$ is countable generated over $F$ by [3, Corollary 5.2]. Consider $U_{i}$, $\mathcal{U}_{i}$, $\widetilde{V}$, $\widetilde{k}$, $\widetilde{\Gamma}$, $\widetilde{s}$, $\widetilde{t}$ given by Theorems 23, 26 for $V$, or the completion $\hat{V}$ of $V$ in case $\dim V=1$. Thus in both case $V$ is Henselian and so $\widetilde{V}$ is Henselian too. Thus we may assume $\widetilde{k}\subset\widetilde{V}$ (see Theorem 26 and Remark 27) and so it is enough to apply Theorem 24. As usual $\widetilde{V}$ is an immediate extension of a valuation ring $)\widetilde{V}_{0}=\widetilde{V}\cap\widetilde{k}(\widetilde{s}(\widetilde{\Gamma}$)), which is a filtered union of localizations of polynomial algebras over $\widetilde{k}$.
Then the map $w^{\prime}:\widetilde{k}\otimes_{F}E\rightarrow\widetilde{V}$ factors through a complete intersection $\widetilde{k}-$algebra and so through a complete intersection $F$-algebra $D$ since $\widetilde{k}/k$ and $k/F$ are separable.
Thus $w$ or the composite map $E\rightarrow V\rightarrow\hat{V}$ when $\dim V=1$ factors through $D$ too because all finite systems of polynomial equations which have a solution in $\tilde{V}$ must have one in $V$, respectively $\hat{V}$ when $\dim V=1$.
Remains the case 2), that is $\dim V=1$. Since $V$ is separable over $F$ we see that $w(H_{E/F})V\not=0$, where $H_{E/F}$ is the ideal defining the non smooth locus of $E$ over $F$ as in Lemma 11. Then $w$ factors also by $D$ by Lemma 20. The proof ends applying [28, Lemma 1.5].
Corollary 7 has the following version when $\Gamma$ is not necessarily free.
Corollary 29.
Let $V$ be a valuation ring containing its residue field $k$ with characteristic $p>0$. Then $V$ is a filtered colimit of some complete intersection $k$-algebras of finite type. Moreover, $V$ is a filtered colimit of some complete intersection ${\bf F}_{p}$-algebras of finite type.

### v12

@@ 1. Filtered colimits of smooth algebras
Remark 20.
If the value group of $V$ is not finitely generated then the above theorem could be false, because the algebraic separable extension $L/k(x)$ could be not dense, as shows [22, Example 9], or [26, Example 8].
@@ 2. Model Theory of valued fields
The conclusion goes from Corollary 7 if we show that $k\subset V$. A lifting of $k$ to $V$ could be done when $V$ is Henselian and char $k=0$ (see [31, Theorem 2.9] but the proof goes in the same way in positive characteristic when $k$ is separably generated over $F$. In particular the lifting could be done when $k/F$ is separable of finite type.
@@
where $\widetilde{K}$ and $\widetilde{\Gamma}$ are the fraction field and the value group of $\widetilde{V}$ respectively.
@@
Let $V$ be a valuation ring containing its residue field $k$ of positive characteristic and $\Gamma$ its value group. Then $V$ is a filtered colimit of complete intersection $k$-algebras. Moreover, $V$ is a filtered colimit of some complete intersection ${\bf F}_{p}$-algebras of finite type.
@@
As $K_{E}$ is a finite type field extension of $k$ we may assume $\Gamma$ is countable generated by [3, Corollary 5.2]. Consider $U_{i}$, $\mathcal{U}_{i}$, $\widetilde{V}$, $\widetilde{k}$, $\widetilde{\Gamma}$, $\widetilde{s}$, $\widetilde{t}$ given by Theorem 23 for $V$. Note that $\widetilde{k}\subset\widetilde{V}$ because $k\subset V$. As usual $\widetilde{V}$ is an immediate extension of a valuation ring $)\widetilde{V}_{0}=\widetilde{V}\cap\widetilde{k}(\widetilde{s}(\widetilde{\Gamma}$)), which is a filtered union of localizations of polynomial algebras over $\widetilde{k}$. Using Theorem 6 we see that $\widetilde{V}$ is a filtered union of its complete intersection $\widetilde{V}_{0}$-subalgebras.
Then the map $w^{\prime}:\widetilde{k}\otimes_{F}E\rightarrow\widetilde{V}$ factors through a complete intersection $\widetilde{k}-$algebra and so the map $E\rightarrow\widetilde{V}$ induced by $w$ factors through a complete intersection $k$-algebra $D$ since $\widetilde{k}/k$ is separable. Thus $w$ factors through $D$ too because all finite systems of polynomial equations which have a solution in $\tilde{V}$ must have one in $V$. The proof ends applying [28, Lemma 1.5].
@@
Let $V$ be a valuation ring with value group $\Gamma$. Then there exists an ultrafilter $\mathcal{U}$ on a set $U$ such that any system of polynomial equations $(g_{i}((X_{j})_{j\in J})_{i\in I}$ with $card\ I\leq$ $card(U)$ in variables $(X_{j})_{j\in J}$ with coefficients in the ultrapower ${\tilde{V}}=\Pi_{\mathcal{U}}V$ has a solution in $\tilde{V}$} if and only if all its finite subsystems have.
@@
Let $V$ be a valuation ring containing a field $F$ of positive characteristic. Assume that $V$ is Henselian and the residue field $k$ of $V$ is a separable extension of $F$. Then $V$ is a filtered colimit of complete intersection $F$-algebras.
@@
Let $\Gamma$ be the value group of $V$. Let $E$ be a finitely generated $F$-subalgebra (in fact finitely presented by [15, Theorem 4]) of $V$ and $w:E\rightarrow V$ the inclusion. We will show that $w$ factors through a complete intersection $V$-algebra. It is enough to replace $V$ by the intersection of $V$ with the fraction field $K_{E}$ of $E$. Thus we may assume that $\Gamma$ is countable generated and $k$ is countable generated over $F$ by [3, Corollary 5.2]. Consider $U_{i}$, $\mathcal{U}_{i}$, $\widetilde{V}$, $\widetilde{k}$, $\widetilde{\Gamma}$, $\widetilde{s}$, $\widetilde{t}$ given by Theorems 23, 26 for $V$. Thus we may assume $\widetilde{k}\subset\widetilde{V}$ (see Theorem 26 and Remark 27) and so it is enough to apply Theorem 24. As usual $\widetilde{V}$ is an immediate extension of a valuation ring $)\widetilde{V}_{0}=\widetilde{V}\cap\widetilde{k}(\widetilde{s}(\widetilde{\Gamma}$)), which is a filtered union of localizations of polynomial algebras over $\widetilde{k}$.
Then the map $w^{\prime}:\widetilde{k}\otimes_{F}E\rightarrow\widetilde{V}$ factors through a complete intersection $\widetilde{k}-$algebra and so the map $E\rightarrow\widetilde{V}$ induced by $w$ factors through a complete intersection $F$-algebra $D$ since $\widetilde{k}/k$ and $k/F$ are separable.
Thus $w$ factors through $D$ too because all finite systems of polynomial equations which have a solution in $\tilde{V}$ must have one in $V$. The proof ends applying [28, Lemma 1.5].